# Model-Agnostic Bounds for Augmented Inverse Probability Weighted Estimators’ Wald-Confidence Interval Coverage in Randomized Controlled Trials

### v1

@@ 6 Discussion
Additionally under Conditions 7 and S1, by Lemma S10, (S7) holds with $\hat{\hat{\mathcal{R}}}(\eta_{k},\lambda_{k,2})$ replaced by $\hat{\hat{\mathcal{R}}}^{\prime}(\eta_{k},\lambda_{k,2},q)$. Taking
$\eta_{k}=\hat{s}_{1}\left[{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}+\hat{s}_{2}\sqrt{2q{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}\log n}\right]\sqrt{2q\frac{|I_{k}|}{n}\log n},$
$\lambda_{k,2}=\frac{|I_{k}|\hat{s}_{3}|z_{\alpha}|M(1-\tau_{\pi})}{n\sigma_{\dagger,a}\tau_{\pi}}\sqrt{2q{\mathbb{E}}\|\hat{Q}_{k,a}-Q_{\#,a}\|_{2}^{2}\log n},$
and $(\lambda_{k,1},\lambda_{k,3})$ the same as above, by Lemma S8 and Theorem S2, (S1) holds with $\mathcal{R}_{k}$ replaced by $\hat{\mathcal{R}}_{k}^{\prime}$.
S4.3 Results for non-cross-fit AIPW procedure
For any function class $\mathcal{G}$} and any $\delta>0$, define ${\mathcal{G}}_{\delta}:=\{f\in{\mathcal{G}}:\|f\|_{2}\leq\delta\}$. Let ${\mathbb{G}}_{n}$ denote the empirical process, so ${\mathbb{G}}_{n}f=\sqrt{n}(P_{n}-P_{*})f$ for any function $f\in L_{2}(P_{*})$, and $\|{\mathbb{G}}_{n}\|_{\mathcal{G}}:=\sup_{f\in{\mathcal{G}}}|{\mathbb{G}}_{n}f|$ for any function class ${\mathcal{G}}\subseteq L_{2}(P_{*})$. Let $\|\cdot\|_{2,\mu}$ denote the $L_{2}(\mu)$ norm for any measure $\mu$. Define ${\mathcal{F}}^{H_{a}}:=\{H_{a}(Q):Q\in{\mathcal{F}}\}$, ${\mathcal{F}}^{{\mathcal{T}}_{a}^{2}}:=\{{\mathcal{T}}_{a}(Q)^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}:Q\in{\mathcal{F}}\}$, where $\mathcal{F}$} is introduced in Condition 9. Recall $F^{H_{a}}$ and $F^{{\mathcal{T}}_{a}^{2}}$ defined above Theorem S3.
@@
Suppose that Conditions 3, 4, and 9 hold. For any $a\in\{0,1\}$, ${\mathcal{F}}^{H_{a}}$ and ${\mathcal{F}}^{{\mathcal{T}}_{a}^{2}}$ have constant envelopes $F^{H_{a}}$ and $F^{{\mathcal{T}}_{a}^{2}}$, respectively. Moreover, for any probability measure $\mu$,
@@
For any $\delta>0$, $\alpha>0$, $q\geq 2$ and $t\geq 1$, under Conditions 3, 4 and 9,
$\max\left\{{\mathbb{E}}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}^{H_{a}}_{\delta F^{H_{a}}}},{\mathbb{E}}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}^{{\mathcal{T}}_{a}^{2}}_{\delta F^{{\mathcal{T}}_{a}^{2}}}}\right\}\leq{\mathcal{C}}F^{H_{a}}J(2\delta,{\mathcal{F}},M)+{\mathcal{C}}\frac{F^{H_{a}}J^{2}(2\delta,{\mathcal{F}},M)}{\delta^{2}\sqrt{n}},$
$\max\left\{{\mathbb{E}}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}^{H_{a}}},{\mathbb{E}}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}^{{\mathcal{T}}_{a}^{2}}}\right\}\leq{\mathcal{C}}F^{H_{a}}J(2,{\mathcal{F}},M),$
@@
$\Pr\left(|{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})|>(1+\alpha){\mathbb{E}}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}^{H_{a}}_{\delta F^{H_{a}}}}+{\mathcal{C}}q\left\{\left(\delta F^{H_{a}}+\frac{F^{H_{a}}}{\sqrt{n}}\right)\sqrt{t}+\frac{F^{H_{a}}t}{\alpha\sqrt{n}}\right\},\|H_{a}(\tilde{Q}_{a})\|_{2}\leq\delta F^{H_{a}}\right),$
$\Pr\Big(|{\mathbb{G}}_{n}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}|>(1+\alpha){\mathbb{E}}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}^{{\mathcal{T}}_{a}^{2}}_{\delta F^{{\mathcal{T}}_{a}^{2}}}}+{\mathcal{C}}q\left\{\left(\delta F^{{\mathcal{T}}_{a}^{2}}+\frac{F^{{\mathcal{T}}_{a}^{2}}}{\sqrt{n}}\right)\sqrt{t}+\frac{F^{{\mathcal{T}}_{a}^{2}}t}{\alpha\sqrt{n}}\right\},$
@@
$\Pr\left(|{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})|>(1+\alpha){\mathbb{E}}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}^{H_{a}}}+{\mathcal{C}}q\left\{F^{H_{a}}\sqrt{t}+\frac{F^{H_{a}}t}{\alpha\sqrt{n}}\right\}\right).$
@@
$\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}_{\delta F^{H_{a}}}^{H_{a}}}\leq(1+\alpha){\mathbb{E}}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}_{\delta F^{H_{a}}}^{H_{a}}}+{\mathcal{C}}q\left\{\left(\delta F^{H_{a}}+\frac{F^{H_{a}}}{\sqrt{n}}\right)\sqrt{t}+\frac{F^{H_{a}}t}{\alpha\sqrt{n}}\right\}.$
@@
${\mathbb{E}}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}_{\delta F^{H_{a}}}^{H_{a}}}\leq{\mathcal{C}}F^{H_{a}}J(\delta,{\mathcal{F}}^{H_{a}},F^{H_{a}})+{\mathcal{C}}\frac{F^{H_{a}}J^{2}(\delta,{\mathcal{F}}^{H_{a}},F^{H_{a}})}{\delta^{2}\sqrt{n}}.$
The desired result follows from Condition 9 and Lemma S11.
The result for $H_{a}(\tilde{Q}_{a})$ without $\|H_{a}(\tilde{Q}_{a})\|_{2}\leq\delta F^{H_{a}}$ is also similar, except that ${\mathcal{F}}^{H_{a}}$ is considered so that $\sigma$ in Theorem 5.2 in Chernozhukov et al., (2014) is taken to be 1, and, by the proof of Theorem 2.5.2 in van der Vaart and Wellner, (1996) (particularly the fourth displayed equation on page 128), ${\mathbb{E}}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}^{H_{a}}}\leq{\mathcal{C}}F^{H_{a}}J(1,{\mathcal{F}}^{H_{a}},F^{H_{a}})$. The bounds for ${\mathcal{F}}_{{\mathcal{T}}_{a}^{2}}$ and ${\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}$ can be shown similarly. ∎
@@
Under Conditions 1, 3, 4 and 9, $|{\mathbb{E}}[\tilde{\psi}_{a}]-\psi_{*,a}|\leq{\mathcal{C}}n^{-1/2}J(1,{\mathcal{F}}^{H_{a}},F^{H_{a}})F^{J_{a}}$,
@@
$\Pr\left(\left|\tilde{\psi}_{a}^{2}-{\mathbb{E}}[\tilde{\psi}_{a}^{2}]\right|>\sum_{j=1}^{5}\eta_{j}\right)$
@@
Since $\tilde{\psi}_{a}=P_{n}{\mathcal{T}}_{a}(\tilde{Q}_{a})$ and $P_{*}{\mathcal{T}}_{a}(Q_{\#,a})=P_{*}{\mathcal{T}}(\tilde{Q}_{a})=\psi_{*,a}$, it holds that $\tilde{\psi}_{a}=\psi_{*,a}+(P_{n}-P_{*}){\mathcal{T}}_{a}(Q_{\#,a})+n^{-1/2}{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})$. Thus, ${\mathbb{E}}[\tilde{\psi}_{a}]=\psi_{*,a}+n^{-1/2}{\mathbb{E}}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]$, and so $|{\mathbb{E}}[\tilde{\psi}_{a}]-\psi_{*,a}|$ can be bounded using Lemma S12.
By Theorem 2.14.1 in van der Vaart and Wellner, (1996), ${\mathbb{E}}[\{{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})\}^{2}]\leq{\mathcal{C}}J(1,{\mathcal{F}}^{H_{a}},F^{H_{a}})^{2}(F^{H_{a}})^{2}$. Thus,
@@
${\mathbb{E}}[\tilde{\psi}_{a}^{2}]$ $=({\mathbb{E}}[\tilde{\psi}_{a}])^{2}+{\mathrm{Var}}(\tilde{\psi}_{a})=\psi_{*,a}^{2}+n^{-1}\{{\mathbb{E}}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]\}^{2}+2\psi_{*,a}n^{-1/2}{\mathbb{E}}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]+{\mathrm{Var}}(\tilde{\psi}_{a}).$
@@
$\Pr\left(\left|\tilde{\psi}_{a}^{2}-{\mathbb{E}}[\tilde{\psi}_{a}^{2}]\right|>\sum_{j=1}^{5}\eta_{j}\right)$
$=\Pr\Bigg(\Bigg|\{(P_{n}-P_{*}){\mathcal{T}}_{a}(Q_{\#,a})\}^{2}+n^{-1}\{{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})\}^{2}$
@@
$\qquad-\left\{n^{-1}\{{\mathbb{E}}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]\}^{2}+2\psi_{*,a}n^{-1/2}{\mathbb{E}}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]+{\mathrm{Var}}(\tilde{\psi}_{a})\right\}\Bigg|>\sum_{j=1}^{5}\eta_{j}\Bigg)$
$\leq\Pr\left(\{(P_{n}-P_{*}){\mathcal{T}}_{a}(Q_{\#,a})\}^{2}>\eta_{1}\right)+\Pr\left(|n^{-1}\{{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})\}^{2}|>\eta_{2}\right)+\Pr\left(|2\psi_{*,a}(P_{n}-P_{*}){\mathcal{T}}_{a}(Q_{\#,a})|>\eta_{3}\right)$
$\quad+\Pr\left(|2\psi_{*,a}n^{-1/2}{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})-2\psi_{*,a}n^{-1/2}{\mathbb{E}}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]-\left\{n^{-1}\{{\mathbb{E}}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]\}^{2}+{\mathrm{Var}}(\tilde{\psi}_{a})\right\}|>\eta_{4}\right)$
$\quad+\Pr\left(|2n^{-1/2}(P_{n}-P_{*}){\mathcal{T}}_{a}(Q_{\#,a})\cdot{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})|>\eta_{5}\right).$
@@
Term 1: Term 1 equals $\Pr\left(|{\mathbb{G}}_{n}{\mathcal{T}}_{a}(Q_{\#,a})|/\sigma_{\#,a}>\sqrt{\eta_{1}n}/\sigma_{\#,a}\right)$. By Berry-Esseen bound and Gaussian tail bound, Term 1 is upper bounded by
@@
Term 2: Term 2 equals $\Pr(|{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})|>\sqrt{\eta_{2}n})$. With the chosen $\eta_{2}$, by Lemma S12, Term 2 is upper bounded by $t_{2}^{-q_{2}/2}$.
Term 3: If $\psi_{*,a}=0$, then Term 3 equals 0. Otherwise, Term 3 equals $\Pr(|{\mathbb{G}}_{n}{\mathcal{T}}_{a}(Q_{\#,a})|>\eta_{3}\sqrt{n}/(2|\psi_{*,a}|))$. By Berry-Esseen bound and Gaussian tail bound, Term 3 is upper bounded by
@@
$\qquad+\left|2\psi_{*,a}n^{-1/2}{\mathbb{E}}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]+n^{-1}\{{\mathbb{E}}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]\}^{2}+{\mathrm{Var}}(\tilde{\psi}_{a})\right|$
$\geq\left|2\psi_{*,a}n^{-1/2}{\mathbb{E}}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]+n^{-1}\{{\mathbb{E}}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]\}^{2}+{\mathrm{Var}}(\tilde{\psi}_{a})\right|.$
@@
$\Pr\left(|2\psi_{*,a}n^{-1/2}{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})|>\eta_{4}-\left|2\psi_{*,a}n^{-1/2}{\mathbb{E}}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]+n^{-1}\{{\mathbb{E}}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]\}^{2}+{\mathrm{Var}}(\tilde{\psi}_{a})\right|\right)$
$=\Pr\left(|{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})|>\sqrt{n}\frac{\eta_{4}-\left|2\psi_{*,a}n^{-1/2}{\mathbb{E}}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]+n^{-1}\{{\mathbb{E}}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]\}^{2}+{\mathrm{Var}}(\tilde{\psi}_{a})\right|}{2|\psi_{*,a}|}\right)$
$\leq\Pr\left(|{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})|>{\mathcal{C}}(1+\alpha_{4})J(2,{\mathcal{F}},M)F^{H_{a}}+{\mathcal{C}}q_{4}\left\{F^{H_{a}}\sqrt{t_{4}}+\frac{F^{H_{a}}t_{4}}{\alpha_{4}\sqrt{n}}\right\}\right)$
@@
Term 5: Since $|{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})|\leq 2F^{H_{a}}$, Term 5 is upper bounded by $\Pr(|{\mathbb{G}}_{n}{\mathcal{T}}_{a}(Q_{\#,a})|/\sigma_{\#,a}>\eta_{5}n/(4F^{H_{a}}\sigma_{\#,a}))$, which, by Berry-Esseen bound and Gaussian tail bound, is further upper bounded by
@@
Suppose that Conditions 15 and 9 hold. Recall the constant $\varrho_{\#,a}$ defined above Theorem S3. For any positive constants $\delta,\alpha_{6},q_{6}\geq 2,t_{6}\geq 1,\eta_{7},\eta_{8}$, with
@@
$\Pr\left(|\tilde{\sigma}_{a}^{2}-{\mathbb{E}}[\tilde{\sigma}_{a}^{2}]|>\sum_{j=1}^{8}\eta_{j},\|{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\|_{2}\leq\delta F^{{\mathcal{T}}_{a}^{2}}\right)$

### v2

@@ 6 Discussion
Additionally under Conditions 7 and 8, by Lemma S10, (S7) holds with $\hat{\hat{\mathcal{R}}}(\eta_{k},\lambda_{k,2})$ replaced by $\hat{\hat{\mathcal{R}}}^{\prime}(\eta_{k},\lambda_{k,2},q)$. Taking
$\eta_{k}=\left\{2[(q+\hat{d}_{1})\log n]^{(\hat{q}_{1}+2)/(2\hat{q}_{1})}\hat{c}_{1}^{-1/\hat{q}_{1}}\hat{r}_{1}(n)\sqrt{\frac{\tau_{\pi}}{1-\tau_{\pi}}\frac{|I_{k}|}{n}}\right\}\vee\left\{\frac{8}{3}M\frac{1-\tau_{\pi}}{\tau_{\pi}}\frac{(q+\hat{d}_{1})\log n}{\sqrt{n}}\right\},$
$\lambda_{k,2}=\left[\frac{(q+\hat{d}_{2})\log n}{\hat{c}_{2}}\right]^{1/\hat{q}_{2}}\frac{|I_{k}||z_{\alpha}|}{n\sigma_{\dagger,a}}\hat{r}_{2}(n),$
and $(\lambda_{k,1},\lambda_{k,3})$ the same as above, so that $\eta^{\prime}(\eta_{k})\geq(q+\hat{d}_{1})\log n$, by Lemma S8 and Theorem S2, (S1) holds with $\mathcal{R}_{k}$ replaced by $\hat{\mathcal{R}}_{k}^{\prime}$.
S5.3 Results for non-cross-fit AIPW procedure
For any function class $\mathcal{G}$} and any $\delta>0$, define ${\mathcal{G}}_{\delta}:=\{f\in{\mathcal{G}}:\|f\|_{2}\leq\delta\}$. Let ${\mathbb{G}}_{n}$ denote the empirical process, so ${\mathbb{G}}_{n}f=\sqrt{n}(P_{n}-P_{*})f$ for any function $f\in L_{2}(P_{*})$, and $\|{\mathbb{G}}_{n}\|_{\mathcal{G}}:=\sup_{f\in{\mathcal{G}}}|{\mathbb{G}}_{n}f|$ for any function class ${\mathcal{G}}\subseteq L_{2}(P_{*})$. Let $\|\cdot\|_{2,\mu}$ denote the $L_{2}(\mu)$ norm for any measure $\mu$. Define ${\mathcal{F}}^{H_{a}}:=\{H_{a}(Q):Q\in{\mathcal{F}}\}$, ${\mathcal{F}}^{{\mathcal{T}}_{a}^{2}}:=\{{\mathcal{T}}_{a}(Q)^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}:Q\in{\mathcal{F}}\}$, where $\mathcal{F}$} is introduced in Condition 10. Recall $F^{H_{a}}$ and $F^{{\mathcal{T}}_{a}^{2}}$ defined above Theorem S3.
@@
Suppose that Conditions 3, 4, and 10 hold. For any $a\in\{0,1\}$, ${\mathcal{F}}^{H_{a}}$ and ${\mathcal{F}}^{{\mathcal{T}}_{a}^{2}}$ have constant envelopes $F^{H_{a}}$ and $F^{{\mathcal{T}}_{a}^{2}}$, respectively. Moreover, for any probability measure $\mu$,
@@
For any $\delta>0$, $\alpha>0$, $q\geq 2$ and $t\geq 1$, under Conditions 3, 4 and 10,
$\max\left\{{E}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}^{H_{a}}_{\delta F^{H_{a}}}},{E}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}^{{\mathcal{T}}_{a}^{2}}_{\delta F^{{\mathcal{T}}_{a}^{2}}}}\right\}\leq{\mathcal{C}}F^{H_{a}}J(2\delta,{\mathcal{F}},M)+{\mathcal{C}}\frac{F^{H_{a}}J^{2}(2\delta,{\mathcal{F}},M)}{\delta^{2}\sqrt{n}},$
$\max\left\{{E}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}^{H_{a}}},{E}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}^{{\mathcal{T}}_{a}^{2}}}\right\}\leq{\mathcal{C}}F^{H_{a}}J(2,{\mathcal{F}},M),$
@@
${\mathrm{pr}}\left(|{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})|>(1+\alpha){E}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}^{H_{a}}_{\delta F^{H_{a}}}}+{\mathcal{C}}q\left\{\left(\delta F^{H_{a}}+\frac{F^{H_{a}}}{\sqrt{n}}\right)\sqrt{t}+\frac{F^{H_{a}}t}{\alpha\sqrt{n}}\right\},\|H_{a}(\tilde{Q}_{a})\|_{2}\leq\delta F^{H_{a}}\right),$
${\mathrm{pr}}\Big(|{\mathbb{G}}_{n}\{{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\}|>(1+\alpha){E}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}^{{\mathcal{T}}_{a}^{2}}_{\delta F^{{\mathcal{T}}_{a}^{2}}}}+{\mathcal{C}}q\left\{\left(\delta F^{{\mathcal{T}}_{a}^{2}}+\frac{F^{{\mathcal{T}}_{a}^{2}}}{\sqrt{n}}\right)\sqrt{t}+\frac{F^{{\mathcal{T}}_{a}^{2}}t}{\alpha\sqrt{n}}\right\},$
@@
${\mathrm{pr}}\left(|{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})|>(1+\alpha){E}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}^{H_{a}}}+{\mathcal{C}}q\left\{F^{H_{a}}\sqrt{t}+\frac{F^{H_{a}}t}{\alpha\sqrt{n}}\right\}\right).$
@@
$\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}_{\delta F^{H_{a}}}^{H_{a}}}\leq(1+\alpha){E}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}_{\delta F^{H_{a}}}^{H_{a}}}+{\mathcal{C}}q\left\{\left(\delta F^{H_{a}}+\frac{F^{H_{a}}}{\sqrt{n}}\right)\sqrt{t}+\frac{F^{H_{a}}t}{\alpha\sqrt{n}}\right\}.$
@@
${E}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}_{\delta F^{H_{a}}}^{H_{a}}}\leq{\mathcal{C}}F^{H_{a}}J(\delta,{\mathcal{F}}^{H_{a}},F^{H_{a}})+{\mathcal{C}}\frac{F^{H_{a}}J^{2}(\delta,{\mathcal{F}}^{H_{a}},F^{H_{a}})}{\delta^{2}\sqrt{n}}.$
The desired result follows from Condition 10 and Lemma S11.
The result for $H_{a}(\tilde{Q}_{a})$ without $\|H_{a}(\tilde{Q}_{a})\|_{2}\leq\delta F^{H_{a}}$ is also similar, except that ${\mathcal{F}}^{H_{a}}$ is considered so that $\sigma$ in Theorem 5.2 in Chernozhukov et al., (2014) is taken to be 1, and, by the proof of Theorem 2.5.2 in van der Vaart and Wellner, (2023) (particularly the fourth displayed equation on page 193), ${E}\|{\mathbb{G}}_{n}\|_{{\mathcal{F}}^{H_{a}}}\leq{\mathcal{C}}F^{H_{a}}J(1,{\mathcal{F}}^{H_{a}},F^{H_{a}})$. The bounds for ${\mathcal{F}}_{{\mathcal{T}}_{a}^{2}}$ and ${\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}$ can be shown similarly. ∎
@@
Under Conditions 1, 3, 4 and 10, $|{E}[\tilde{\psi}_{a}]-\psi_{*,a}|\leq{\mathcal{C}}n^{-1/2}J(1,{\mathcal{F}}^{H_{a}},F^{H_{a}})F^{J_{a}}$,
@@
${\mathrm{pr}}\left(\left|\tilde{\psi}_{a}^{2}-{E}[\tilde{\psi}_{a}^{2}]\right|>\sum_{j=1}^{5}\eta_{j}\right)$
@@
Since $\tilde{\psi}_{a}=P_{n}{\mathcal{T}}_{a}(\tilde{Q}_{a})$ and $P_{*}{\mathcal{T}}_{a}(Q_{\#,a})=P_{*}{\mathcal{T}}(\tilde{Q}_{a})=\psi_{*,a}$, it holds that $\tilde{\psi}_{a}=\psi_{*,a}+(P_{n}-P_{*}){\mathcal{T}}_{a}(Q_{\#,a})+n^{-1/2}{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})$. Thus, ${E}[\tilde{\psi}_{a}]=\psi_{*,a}+n^{-1/2}{E}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]$, and so $|{E}[\tilde{\psi}_{a}]-\psi_{*,a}|$ can be bounded using Lemma S12.
By Theorem 2.14.1 in van der Vaart and Wellner, (2023), ${E}[\{{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})\}^{2}]\leq{\mathcal{C}}J(1,{\mathcal{F}}^{H_{a}},F^{H_{a}})^{2}(F^{H_{a}})^{2}$. Thus,
@@
${E}[\tilde{\psi}_{a}^{2}]$ $=({E}[\tilde{\psi}_{a}])^{2}+{\mathrm{var}}(\tilde{\psi}_{a})=\psi_{*,a}^{2}+n^{-1}\{{E}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]\}^{2}+2\psi_{*,a}n^{-1/2}{E}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]+{\mathrm{var}}(\tilde{\psi}_{a}).$
@@
${\mathrm{pr}}\left(\left|\tilde{\psi}_{a}^{2}-{E}[\tilde{\psi}_{a}^{2}]\right|>\sum_{j=1}^{5}\eta_{j}\right)$
$={\mathrm{pr}}\Bigg(\Bigg|\{(P_{n}-P_{*}){\mathcal{T}}_{a}(Q_{\#,a})\}^{2}+n^{-1}\{{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})\}^{2}$
@@
$\qquad-\left\{n^{-1}\{{E}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]\}^{2}+2\psi_{*,a}n^{-1/2}{E}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]+{\mathrm{var}}(\tilde{\psi}_{a})\right\}\Bigg|>\sum_{j=1}^{5}\eta_{j}\Bigg)$
$\leq{\mathrm{pr}}\left(\{(P_{n}-P_{*}){\mathcal{T}}_{a}(Q_{\#,a})\}^{2}>\eta_{1}\right)+{\mathrm{pr}}\left(|n^{-1}\{{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})\}^{2}|>\eta_{2}\right)+{\mathrm{pr}}\left(|2\psi_{*,a}(P_{n}-P_{*}){\mathcal{T}}_{a}(Q_{\#,a})|>\eta_{3}\right)$
$\quad+{\mathrm{pr}}\left(|2\psi_{*,a}n^{-1/2}{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})-2\psi_{*,a}n^{-1/2}{E}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]-\left\{n^{-1}\{{E}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]\}^{2}+{\mathrm{var}}(\tilde{\psi}_{a})\right\}|>\eta_{4}\right)$
$\quad+{\mathrm{pr}}\left(|2n^{-1/2}(P_{n}-P_{*}){\mathcal{T}}_{a}(Q_{\#,a})\cdot{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})|>\eta_{5}\right).$
@@
Term 1: Term 1 equals ${\mathrm{pr}}\left(|{\mathbb{G}}_{n}{\mathcal{T}}_{a}(Q_{\#,a})|/\sigma_{\#,a}>\sqrt{\eta_{1}n}/\sigma_{\#,a}\right)$. By Berry-Esseen bound and Gaussian tail bound, Term 1 is upper bounded by
@@
Term 2: Term 2 equals ${\mathrm{pr}}(|{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})|>\sqrt{\eta_{2}n})$. With the chosen $\eta_{2}$, by Lemma S12, Term 2 is upper bounded by $t_{2}^{-q_{2}/2}$.
Term 3: If $\psi_{*,a}=0$, then Term 3 equals 0. Otherwise, Term 3 equals ${\mathrm{pr}}(|{\mathbb{G}}_{n}{\mathcal{T}}_{a}(Q_{\#,a})|>\eta_{3}\sqrt{n}/(2|\psi_{*,a}|))$. By Berry-Esseen bound and Gaussian tail bound, Term 3 is upper bounded by
@@
$\qquad+\left|2\psi_{*,a}n^{-1/2}{E}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]+n^{-1}\{{E}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]\}^{2}+{\mathrm{var}}(\tilde{\psi}_{a})\right|$
$\geq\left|2\psi_{*,a}n^{-1/2}{E}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]+n^{-1}\{{E}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]\}^{2}+{\mathrm{var}}(\tilde{\psi}_{a})\right|.$
@@
${\mathrm{pr}}\left(|2\psi_{*,a}n^{-1/2}{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})|>\eta_{4}-\left|2\psi_{*,a}n^{-1/2}{E}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]+n^{-1}\{{E}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]\}^{2}+{\mathrm{var}}(\tilde{\psi}_{a})\right|\right)$
$={\mathrm{pr}}\left(|{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})|>\sqrt{n}\frac{\eta_{4}-\left|2\psi_{*,a}n^{-1/2}{E}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]+n^{-1}\{{E}[{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})]\}^{2}+{\mathrm{var}}(\tilde{\psi}_{a})\right|}{2|\psi_{*,a}|}\right)$
$\leq{\mathrm{pr}}\left(|{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})|>{\mathcal{C}}(1+\alpha_{4})J(2,{\mathcal{F}},M)F^{H_{a}}+{\mathcal{C}}q_{4}\left\{F^{H_{a}}\sqrt{t_{4}}+\frac{F^{H_{a}}t_{4}}{\alpha_{4}\sqrt{n}}\right\}\right)$
@@
Term 5: Since $|{\mathbb{G}}_{n}H_{a}(\tilde{Q}_{a})|\leq 2F^{H_{a}}$, Term 5 is upper bounded by ${\mathrm{pr}}(|{\mathbb{G}}_{n}{\mathcal{T}}_{a}(Q_{\#,a})|/\sigma_{\#,a}>\eta_{5}n/(4F^{H_{a}}\sigma_{\#,a}))$, which, by Berry-Esseen bound and Gaussian tail bound, is further upper bounded by
@@
Suppose that Conditions 15 and 10 hold. Recall the constant $\varrho_{\#,a}$ defined above Theorem S3. For any positive constants $\delta,\alpha_{6},q_{6}\geq 2,t_{6}\geq 1,\eta_{7},\eta_{8}$, with
@@
${\mathrm{pr}}\left(|\tilde{\sigma}_{a}^{2}-{E}[\tilde{\sigma}_{a}^{2}]|>\sum_{j=1}^{8}\eta_{j},\|{\mathcal{T}}_{a}(\tilde{Q}_{a})^{2}-{\mathcal{T}}_{a}(Q_{\#,a})^{2}\|_{2}\leq\delta F^{{\mathcal{T}}_{a}^{2}}\right)$